OpenStudy (anonymous):

Solve the quadratic equation. (x - 5)2 = 9 A) x = 5 or 3 B) x = 8 or 2 C) x = -5 or 3 D) x = -8 or 2

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ (x - 5)^2 = 9}\) \(\normalsize\color{black}{ (x - 5)^2 = (±3)^2}\) \(\normalsize\color{black}{ x - 5 = ±3}\) \(\normalsize\color{black}{ x = 5±3}\) \(\normalsize\color{black}{ x =?}\)

OpenStudy (anonymous):

Thanks so much! :)